# A method to derive material-specific spin-bath model descriptions of materials displaying prevalent spin physics (for simulation on NISQ devices)

### v1

@@ I Introduction
The theoretical study of the physics of many-body quantum system has remained notoriously difficult on conventional computers not least due to the intractable exponential growth of the corresponding Hilbert spaces. This holds true already for simple model Hamiltonians, but becomes even more problematic for the study of real materials due to the vast number of degrees of freedom and interactions. To investigate certain aspects of these materials it is therefore imperative to either utilize significant approximations or to obtain a much more concise Hamiltonian focused on accurately describing this aspect. In this article, we concentrate on the description of the spin physics of actual materials. The spin physics is relevant for the understanding of the material’s magnetic properties as well as their experimental identification with spin resonance spectroscopy techniques [1]. The quantum simulation of spin system Hamiltonians is also perfectly suited as an application of quantum computing in general [2] and analogue quantum computing in particular.
We present a method for the derivation of the effective Hamiltonian description for the relevant spin degrees of freedom of an actual material embedded in an environment of fermions. The first part of the method entails the identification of potential relevant orbital spin degrees of freedom. We propose a new metric for the classification of orbital spin degrees of freedom and present a systematic way to determine the orbital basis that contains the best realizations of these spins, see also  [3]. Given the presence of relevant spins, we proceed with the second part of the method, a Schrieffer-Wolff transformation procedure [4, 5] aimed at decoupling the subspace of the Hilbert space containing these spin degrees of freedom from the remaining high-energy subspaces. Many different Schrieffer-Wolff and related unitary transformation schemes have been proposed, comprising perturbative [4, 6], variational [7], and continuous [8, 9, 10] variants. So far, the application of these methods has been limited to concise model Hamiltonians. We present an extended version of the perturbative Schrieffer-Wolff transformation that can be applied to generic Hamiltonian descriptions of materials, given the presence of relevant low-energy spin physics in the material. We demonstrate the accuracy of the effective spin-bath model Hamiltonians resulting from our Schrieffer-Wolff transformation for the description of the low-energy physics both for model Hamiltonians with established results [4, 11] as well as an actual material. We further discuss the extent to which the quality of the effective spin-bath model Hamiltonian derived with our proposed method can be predicted.
@@ II.1 Spin-like orbitals
We consider an orbital $\phi_{i}$ as being spin-like only if the electron density $n_{i}$ contained in it is strictly equal to one. This requires that the average electron density in the orbitals $\phi_{i}$ satisfies $\langle n_{i}\rangle=1$. It furthermore requires that the fluctuations around this average electron density satisfy $\delta n_{i}\rightarrow 0$. Negligible fluctuations around the average electron density imply that electron density of the orbital does not affect the low-energy dynamics of the system and vice versa. Just an average electron density $\langle n_{i}\rangle=1$ places no restrictions on the orientation and dynamics of the electron spin in the orbital $\phi_{i}$. When both requirements are simultaneously met by the states in the low-energy Hilbert space, the dynamics of the electron density in the orbital $\phi_{i}$, often referred to as the charge degree of freedom, becomes superfluous to the description of the dynamics of the system. The electron density in the orbital $\phi_{i}$ consequently couples to the remainder of the system exclusively via its spin degree of freedom. It is then sufficient to represent the degrees of freedom of the electron density contained in the orbital $\phi_{i}$ as a pure spin degree of freedom and to employ the associated spin operator algebra. The local Hilbert space $\mathcal{H}_{i}$ of a spin degree of freedom is half the size of the local Hilbert space of a fermionic orbital and is furthermore naturally represented on a Qubit. We therefore aim to determine each spin-like basis orbital or linear combinations thereof meeting both stated requirements, so that they can be represented as spins.
@@ II.3 Local parity optimization of the orbital basis
with the same rotation being performed for the Hermitian conjugates of the operators. From the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ we can compute the local parity of an orbital $\phi_{q}$, which results from the linear combination of orbitals $\phi_{i}$ and $\phi_{j}$, as $\langle P_{q}\rangle_{0}(\theta)$. The local parity $\langle P_{q}\rangle_{0}(\theta)$ is an analytic, $2\pi$-periodic function of the rotation angle $\theta$. We find the extremal points $\theta_{n}$ of the function $\langle P_{q}\rangle_{0}(\theta)$ in the domain $\theta\in[0,2\pi)$ from
@@ III.3 Single impurity Anderson model
For values of the interaction strength $U/t\leq 4$, we observe several pronounced energy differences between the SIAM ground states and the ground states of the transformed and effective Hamiltonians. In this weak interaction regime, the chemical potential of the impurity lies within the bandwith $D=4t$ of the tight-binding chain representing the remainder of the system. There is then no energy gap between the subspace of the Hilbert space where the impurity is singly occupied and the rest of the Hilbert space. In the absence of an energy gap, limiting the series expansion (18) to first order in $S$ is not a justified approximation. For values of the interaction strength $U/t>4$ we observe a decrease of the energy discrepancy $\Delta E_{0}$ of the ground state energies of the SIAM Hamiltonian and the transformed and effective Hamiltonians respectively as a power-law $(U/t)^{-\alpha}$ of the interaction strength with an exponent $\alpha>2$. The energy discrepancy quickly becomes small with respect to the energy scale $t$ and we find $\Delta E_{0}/t\rightarrow 0$ for $U/t\rightarrow\infty$. For the ground state local parity on the impurity site we observe $7\times 10^{-2}>(\langle P_{d}\rangle_{0}+1)>3\times 10^{-2}$ in the regime $U/t\leq 4$. For larger values of the interaction strength $U/t>4$ we find that the local parity of the impurity site approximates a polynomial function $(\langle P_{d}\rangle_{0}+1)(U/t)\propto(U/t)^{-\beta}$ with the exponent $\beta\approx\alpha-1$. The local parity of the impurity site appears to be a good indicator of quality of the approximation (19) if $\tilde{H}$ is supposed to become block-diagonal in the subspaces of the Hilbert space with different particle number $n_{d}$ on the impurity site. This supports our claim that a local parity $(\langle P_{d}\rangle_{0}+1)\rightarrow 0$ signals that the corresponding site can be represented as a spin degree of freedom and that $\mathcal{P}\tilde{H}\mathcal{P}$ is then a good effective Hamiltonian to describe the low-energy spin dynamics of the system. The energy discrepancy $\Delta E_{0}$ is identical for both $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$. In the case of the SIAM Hamiltonian the Schrieffer-Wolff transformation yields a Hamiltonian $\tilde{H}$ that does no longer couple the distinct subspaces characterized by $n_{d}=1$ and $n_{d}\neq 1$. We find that our Schrieffer-Wolff transformation method recovers the analytically known results and that the ground state of effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ approaches the true ground state as $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-1}$.
@@ III.5 Molecular Chromium bromide
We initially perform a series of electronic structure calculations using the def2-QZVP basis set to generate the information necessary for a search of the active space of orbital basis of the system. We have used the Active Space Finder (ASF), [27] an open-source to assist the user in the selection of an active space. As expected, five Chromium $d$-orbitals and $p$-orbitals from the outer Shell of Bromide are suggested by the ASF and the active space contains $N_{\phi_{i}}=9$ basis orbitals. We have performed a CASSCF calculation using the previously identified orbitals as the active space to compute the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ for the CASSCF ground state. From this CASSCF calculation we have also obtained the one-electron $t^{\sigma\sigma^{\prime}}_{ij}$ and two-electron integrals $V^{\sigma\sigma^{\prime}}_{ijkl}$ in the active space. We can write the Hamiltonian description of the active space of chromium bromide as
@@
where $c^{\dagger}_{i\sigma}$ creates an electron with spin $\sigma$ in the orbital $\phi_{i}$ of the active space. Using the reduced density matrices we have performed a parity optimization of the basis orbitals of the active space. Here, we first ran a parity minimization procedure of the entire basis and identified $N_{\phi_{q}}=3$ spin-like orbital with respective parities $\langle P_{q}\rangle_{0}\leq-9.94\times 10^{-1}$. In a subsequent optimization run have maximized the local parities of the remaining basis orbitals. The local parities of the initial basis orbitals $\phi_{i}$ and the optimized basis orbitals $\phi_{q}$ are shown together with the average electron density $\langle n_{q}\rangle_{0}$ in Fig. 7.
@@ IV Conclusion
Effective model Hamiltonians provide enormous value for the investigation of individual aspects of complex systems. Here, we have presented a procedure for the automatic derivation of model Hamiltonians of coupled spin degrees of freedom embedded in a fermionic environment. These Hamiltonians provide a means to enhance the study of the dominant low temperature spin physics of complex real materials. We have shown that to investigate the spin physics of such materials properly, one first needs to operate within a suitable single particle orbital basis, where a subset of the basis orbitals become realizations of spin degrees of freedom at low energies. We have introduced the local parity as a sensible metric for the spin-like behavior of electron densities in the basis orbitals. We have proven that the natural orbital basis does not generally correspond to the basis that features the spin-like orbitals. We have provided an example toy model for which the natural orbital basis is significantly worse than the initial basis for the study of spin physics. We also studied a well-known singlet-diradical molecule, namely closed configuration para-benzyne, and demonstrated that the two well-known non-trivial spin-like orbitals are correctly determined by our local parity optimization procedure. We find that the optimization of the local parity yields an orbital basis that typically results in a many particle basis of the Hilbert space with small coupling and significant energy gap between the different subspaces. Our extended Schrieffer-Wolff transformation method allows to also transform generic Hamiltonians with complicated interaction terms beyond the density-density interactions of the toy models that the Schrieffer-Wolff transformation has so far mostly been applied to. There is also no need for the block-diagonal part $H_{0}$ to be easily diagonalizable, because no calculation of the spectrum of $H_{0}$ is required in Schrieffer-Wolff transformation our method. We have shown that our Schrieffer-Wolff transformation method reproduces established results for the well-known single impurity Anderson model and the disordered Fermi-Hubbard model Hamiltonians. For these models we corroborate that the local parity of the spin-like orbitals is a good predictor for the quality of the effective spin-bath Hamiltonian as the description of the low energy spin dynamics. Lastly, we have applied the full spin-bath Hamiltonian derivation procedure to molecular chromium bromide. Here, we have shown that the derived effective spin-bath models are good descriptions of the system’s low-energy spin excitations. Future work will address more complex materials, where the Schrieffer-Wolff transformation becomes essential to obtain good effective Hamiltonian descriptions for the low-energy spin dynamics of the material.

### v2

@@ I Introduction
The theoretical study of the physics of many-body quantum system has remained notoriously difficult on conventional computers not least due to the intractable exponential growth of the corresponding Hilbert spaces. This holds true already for simple model Hamiltonians, but becomes even more problematic for the study of real materials due to the vast number of degrees of freedom and interactions. To investigate certain aspects of these materials it is therefore imperative to either utilize significant approximations or to obtain a much more concise Hamiltonian focused on accurately describing this aspect. In this article, we concentrate on the description of the spin physics of actual materials. The spin physics is relevant for the understanding of the material’s magnetic properties [1] as well as their experimental identification with spin resonance spectroscopy techniques [2]. The quantum simulation of spin system Hamiltonians is also perfectly suited as an application of quantum computing in general [3] and analogue quantum computing in particular [4, 5, 6].
We present a method for the derivation of the effective Hamiltonian description for the relevant spin degrees of freedom of an actual material embedded in an environment of fermions. The first part of the method entails the identification of potential relevant orbital spin degrees of freedom. We propose a new metric for the classification of orbital spin degrees of freedom and present a systematic way to determine the orbital basis that contains the best realizations of these spins, see also  [7]. Given the presence of relevant spins, we proceed with the second part of the method, a Schrieffer-Wolff transformation procedure [8, 9] aimed at decoupling the subspace of the Hilbert space containing these spin degrees of freedom from the remaining high-energy subspaces. Many different Schrieffer-Wolff and related unitary transformation schemes have been proposed, comprising perturbative [8, 10, 11], variational [12], and continuous [13, 14, 15] variants. So far, the application of these methods has been limited to concise model Hamiltonians. We present an extended version of the perturbative Schrieffer-Wolff transformation that can be applied to generic Hamiltonian descriptions of materials, given the presence of relevant low-energy spin physics in the material. We demonstrate the accuracy of the effective spin-bath model Hamiltonians resulting from our Schrieffer-Wolff transformation for the description of the low-energy physics both for model Hamiltonians with established results [8, 16] as well as an actual material. We further discuss the extent to which the quality of the effective spin-bath model Hamiltonian derived with our proposed method can be predicted.
@@ II.1 Spin-like orbitals
We consider an orbital $\phi_{i}$ as being spin-like only if the electron density $n_{i}$ contained in it is strictly equal to one. This requires that the average electron density in the orbitals $\phi_{i}$ satisfies $\langle n_{i}\rangle=1$. It furthermore requires that the fluctuations around this average electron density satisfy $\delta n_{i}\rightarrow 0$. Negligible fluctuations around the average electron density imply that the electron density of the orbital does not affect the low-energy dynamics of the system and vice versa. Just an average electron density $\langle n_{i}\rangle=1$ places no restrictions on the orientation and dynamics of the electron spin in the orbital $\phi_{i}$. When both requirements are simultaneously met by the states in the low-energy Hilbert space, the dynamics of the electron density in the orbital $\phi_{i}$, often referred to as the charge degree of freedom, becomes superfluous to the description of the dynamics of the system. The electron density in the orbital $\phi_{i}$ consequently couples to the remainder of the system exclusively via its spin degree of freedom. It is then sufficient to represent the degrees of freedom of the electron density contained in the orbital $\phi_{i}$ as a pure spin degree of freedom and to employ the associated spin operator algebra. The local Hilbert space $\mathcal{H}_{i}$ of a spin degree of freedom is half the size of the local Hilbert space of a fermionic orbital and is furthermore naturally represented on a Qubit. We therefore aim to determine each spin-like basis orbital or linear combinations thereof meeting both stated requirements, so that they can be represented as spins.
@@ II.3 Local parity optimization of the orbital basis
with the same rotation being performed for the Hermitian conjugates of the operators. From the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ we can compute the local parity of an orbital $\phi_{q}$, which results from the linear combination of orbitals $\phi_{i}$ and $\phi_{j}$, as $\langle P_{q}\rangle_{0}(\theta)$, where $\langle\dots\rangle_{0}$ denotes the expectation value with respect to the ground state. The local parity $\langle P_{q}\rangle_{0}(\theta)$ is an analytic, $2\pi$-periodic function of the rotation angle $\theta$. We find the extremal points $\theta_{n}$ of the function $\langle P_{q}\rangle_{0}(\theta)$ in the domain $\theta\in[0,2\pi)$ from
@@ III.3 Single impurity Anderson model
For values of the interaction strength $U/t\leq 4$, we observe a pronounced energy difference between the SIAM ground state and the respective ground states of the transformed and effective Hamiltonians. In this weak interaction regime, the chemical potential of the impurity lies within the bandwith $D=4t$ of the tight-binding chain representing the remainder of the system. There is then no energy gap between the subspace of the Hilbert space where the impurity is singly occupied and the rest of the Hilbert space. In the absence of an energy gap, limiting the series expansion (18) to first order in $S$ is not a justified approximation. For values of the interaction strength $U/t>4$ we observe a decrease of the energy discrepancy $\Delta E_{0}$ of the ground state energies of the SIAM Hamiltonian and the transformed and effective Hamiltonians respectively as a power-law $(U/t)^{-\alpha}$ of the interaction strength with an exponent $\alpha>2$. The energy discrepancy quickly becomes small with respect to the energy scale $t$ and we find $\Delta E_{0}/t\rightarrow 0$ for $U/t\rightarrow\infty$. For the ground state local parity on the impurity site we observe $7\times 10^{-2}>(\langle P_{d}\rangle_{0}+1)>3\times 10^{-2}$ in the regime $U/t\leq 4$. For larger values of the interaction strength $U/t>4$ we find that the local parity of the impurity site approximates a polynomial function $(\langle P_{d}\rangle_{0}+1)(U/t)\propto(U/t)^{-\beta}$ with the exponent $\beta\approx\alpha-1$. The local parity of the impurity site appears to be a good indicator of quality of the approximation (19) if $\tilde{H}$ is supposed to become block-diagonal in the subspaces of the Hilbert space with different particle number $n_{d}$ on the impurity site. This supports our claim that a local parity $(\langle P_{d}\rangle_{0}+1)\rightarrow 0$ signals that the corresponding site can be represented as a spin degree of freedom and that $\mathcal{P}\tilde{H}\mathcal{P}$ is then a good effective Hamiltonian to describe the low-energy spin dynamics of the system. The energy discrepancy $\Delta E_{0}$ is identical for both $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$. In the case of the SIAM Hamiltonian the Schrieffer-Wolff transformation yields a Hamiltonian $\tilde{H}$ that does no longer couple the distinct subspaces characterized by $n_{d}=1$ and $n_{d}\neq 1$. We find that our Schrieffer-Wolff transformation method recovers the analytically known results and that the ground state of effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ approaches the true ground state as $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-1}$.
@@ III.5 Molecular Chromium bromide
We initially perform a series of electronic structure calculations using the def2-QZVP basis set to generate the information necessary for a search of the active space of orbital basis of the system. We have used the Active Space Finder (ASF), [32] an open source package to assist the user in the selection of an active space. As expected, five Chromium $d$-orbitals and $p$-orbitals from the outer Shell of Bromide are suggested by the ASF and the active space contains $N_{\phi_{i}}=9$ basis orbitals. We have performed a CASSCF calculation using the previously identified orbitals as the active space to compute the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ for the CASSCF ground state. From this CASSCF calculation we have also obtained the one-electron $t^{\sigma\sigma^{\prime}}_{ij}$ and two-electron integrals $V^{\sigma\sigma^{\prime}}_{ijkl}$ in the active space. We can write the Hamiltonian description of the active space of chromium bromide as
@@
where $c^{\dagger}_{i\sigma}$ creates an electron with spin $\sigma$ in the orbital $\phi_{i}$ of the active space. Using the reduced density matrices we have performed a parity optimization of the basis orbitals of the active space. Here, we first ran a parity minimization procedure of the entire basis and identified $N_{\phi_{q}}=3$ spin-like orbitals with respective parities $\langle P_{q}\rangle_{0}\leq-9.94\times 10^{-1}$. In a subsequent optimization run have maximized the local parities of the remaining basis orbitals. The local parities of the initial basis orbitals $\phi_{i}$ and the optimized basis orbitals $\phi_{q}$ are shown together with the average electron density $\langle n_{q}\rangle_{0}$ in Fig. 7.
@@ IV Conclusion
Effective model Hamiltonians provide enormous value for the investigation of individual aspects of complex systems. Here, we have presented a procedure for the automatic derivation of model Hamiltonians of coupled spin degrees of freedom embedded in a fermionic environment. These Hamiltonians provide a means to enhance the study of the dominant low temperature spin physics of complex real materials. We have shown that to investigate the spin physics of such materials properly, one first needs to operate within a suitable single particle orbital basis, where a subset of the basis orbitals become realizations of spin degrees of freedom at low energies. We introduced the local parity as a sensible metric for the spin-like behavior of electron densities in the basis orbitals. We have proven that the natural orbital basis does not generally correspond to the basis that features the spin-like orbitals. We have provided an example toy model for which the natural orbital basis is significantly worse than the initial basis for the study of spin physics. We also studied a well-known singlet-diradical molecule, namely closed configuration para-benzyne, and demonstrated that the two well-known non-trivial spin-like orbitals are correctly determined by our local parity optimization procedure. We find that the optimization of the local parity yields an orbital basis that typically results in a many particle basis of the Hilbert space with small coupling and significant energy gap between the different subspaces. Our extended Schrieffer-Wolff transformation method allows to also transform generic Hamiltonians with complicated interaction terms beyond the density-density interactions of the toy models that the Schrieffer-Wolff transformation has so far mostly been applied to. There is also no need for the block-diagonal part $H_{0}$ to be easily diagonalizable, because no calculation of the spectrum of $H_{0}$ is required in Schrieffer-Wolff transformation our method. We have shown that our Schrieffer-Wolff transformation method reproduces established results for the well-known single impurity Anderson model and the disordered Fermi-Hubbard model Hamiltonians. For these models we corroborate that the local parity of the spin-like orbitals is a good predictor for the quality of the effective spin-bath Hamiltonian as the description of the low energy spin dynamics. Lastly, we have applied the full spin-bath Hamiltonian derivation procedure to molecular chromium bromide. Here, we have shown that the derived effective spin-bath models are good descriptions of the system’s low-energy spin excitations. Future work will address more complex materials, where the Schrieffer-Wolff transformation becomes essential to obtain good effective Hamiltonian descriptions for the low-energy spin dynamics of the material.
Data availability
The numerical data can be obtained from the authors on request. The spin finding procedure and the implemented Schrieffer Wolff transformation can be tested on https://cloud.quantumsimulations.de/ using the HQS Spin Mapper module.